# SolarGPT-QA: A Domain-Adaptive Large Language Model for Educational Question Answering in Space Weather and Heliophysics

### v1

@@ Abstract
Solar activity, including solar flares, coronal mass ejections (CMEs), and geomagnetic storms, can significantly impact satellites, aviation, power grids, data centers, and space missions. Extreme solar events can cause substantial economic damage if not predicted in advance, highlighting the importance of accurate forecasting and effective education in space science. Although large language models (LLMs) perform well on general tasks, they often lack domain-specific knowledge and pedagogical capability to clearly explain complex space science concepts.
We introduce SolarGPT-QA, a question answering system based on a domain-adapted large language model built on the LLaMA-3 base model. The model is trained using scientific literature and large-scale question-answer data generated with GPT-4 and refined using Grok-3 in a student-friendly storytelling style. Human pairwise evaluations show that SolarGPT-QA outperforms general-purpose models in zero-shot settings and achieves competitive performance compared to instruction-tuned models for educational explanations in space weather and heliophysics. A small pilot student comprehension study further suggests improved clarity and accessibility of the generated explanations. Ablation experiments indicate that combining domain-adaptive pretraining with pedagogical fine-tuning is important for balancing scientific accuracy and educational effectiveness. This work represents an initial step toward a broader SolarGPT framework for space science education and forecasting.
@@ I Introduction
Heliophysics provides a system-level understanding of the Sun and its influence throughout the heliosphere [34]. Within this broader heliophysical framework, space weather science investigates the complex and dynamic interactions between the Sun, the solar wind, interplanetary space, and planetary environments, and examines how solar activity influences Earth’s space environment and technological systems [45]. Solar phenomena such as solar flares (sudden bursts of electromagnetic radiation from the Sun) [45, 44], coronal mass ejections (CMEs) (large eruptions of magnetized plasma from the solar corona) [54], and solar energetic particles (SEPs) (high-energy charged particles accelerated by flares and CMEs) [42, 2] can trigger geomagnetic storms [19], which disrupt satellite operations, data centers, communication systems, and power grids, and pose significant radiation risks to astronauts and high-altitude aviation. For instance, the ”Carrington Event” in 1859 [33] was the most intense geomagnetic storm on record triggered by an extreme solar flare and CME, the Quebec blackout in 1989 [14] showed the vulnerability of modern technological system to space weather events.
Such disruptions have resulted in economic losses of billions of dollars and, in extreme cases, pose risks to human safety and critical infrastructure [5, 2]. As modern society becomes increasingly dependent on space-based technologies—such as large satellite constellations (e.g., Starlink) [38] and as private space enterprises including SpaceX and Blue Origin pursue future space-based data centers and exploration missions [52], understanding space weather has become a critical scientific, technological, and educational priority, beginning as early as the high-school level.
@@
Large language models have become powerful tools for understanding and explaining complex information, with growing impact across many scientific fields [3]. They can answer questions, summarize research, and support learning. However, their use in space weather and heliophysics is still limited, largely because these areas rely on highly specialized literature and concepts that are rarely covered in general training data. One key challenge is most space weather knowledge lives in technical papers and mission reports, which means general models often miss key terms, cause–effect links, or physical limits. Another issue is teaching space science particularly to K–12 students or early undergraduates requires clear language, analogies, and storytelling approach. General models are not built to balance accuracy that is tailored to the appropriate age.
@@
To fill this gap, we introduce SolarGPT-QA, a domain-adapted model designed for educational question answering in space weather and heliophysics for K–12 learners. The model is trained on a carefully curated corpus of high-impact, peer-reviewed research articles published from 2010 onward, aligned with the Solar Dynamics Observatory (SDO) era of modern flare and CME research. These sources are further refined for relevance to solar eruptions and used to generate a large collection of synthetic question–answer pairs. Rather than targeting expert-level analysis, SolarGPT-QA focuses on helping students grasp core concepts through clear, accessible explanations to questions such as how solar flares form or why the Sun can disrupt technology on Earth.
@@ II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [18]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [28, 37, 20]. Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [53, 57], but their operational use is limited by computational cost and incomplete boundary conditions [56].
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [21] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters, employing decision trees and sequence models to model temporal dependencies [31].
Progress in this area has been supported by the development of benchmark datasets. The SWAN-SF multivariate time series dataset, integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [1]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [16, 22].
Many efforts have used deep learning on solar imagery, including magnetograms and extreme ultraviolet observations, to capture spatial precursors to flares and CMEs, while multimodal fusion of time series and images further enhances forecasting accuracy [36, 58, 23]. Despite these advances, most operational systems predict flares, CMEs, and SEPs independently and provide limited explanatory context, functioning largely as black boxes [13, 39, 40]. These limitations motivate unified, multimodal, and explainable forecasting frameworks that build on prior advances in time series modeling, dataset development, and interpretability.
@@ II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve best performance across wide range of the tasks and model scales [6, 12, 50, 51, 9, 8, 7]. There has been subsequent work and exploring for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [41, 60, 49, 43, 59]. DAPT and finetuning are common techniques to improve the LLM for specific dataset or specific domain. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [26, 24, 17]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [55, 35, 47, 11].
@@ III SolarGPT-QA
SolarGPT-QA is a domain-adapted large language model designed specifically for educational question answering in space weather and heliophysics, with an emphasis on K–12–friendly explanations. The methodology consists of four main stages: (1) domain corpus construction and preprocessing, (2) domain-adaptive pretraining (DAPT) using LoRA, (3) educational question–answer fine-tuning, and (4) inference-time constraints to promote clarity and reduce hallucinations. An overview of the system architecture is shown in Fig. 2. Let $\mathcal{D}_{H}$ denote a corpus of heliophysics literature and $\mathcal{D}_{QA}$ denote a supervised question–answer dataset designed for K–12 learners. Our objective is to learn a conditional language model
@@
where $x$ is a heliophysics question and $y$ is an age-appropriate explanatory answer that preserves scientific correctness while prioritizing conceptual clarity.
@@ III-A Domain Corpus Construction and Normalization
We curate a heliophysics corpus $\mathcal{D}_{H}=\{d_{i}\}_{i=1}^{N}$ from peer-reviewed literature published between 2010 and 2024, aligned with the operational era of the Solar Dynamics Observatory (SDO). Each document $d_{i}$ is extracted from PDF format and transformed using a deterministic preprocessing function:
@@ III-B Domain-Adaptive Pretraining with Low-Rank Adaptation
The overall training workflow is summarized in Algorithm 1. Only the low-rank matrices $A$ and $B$ are trainable, while the base model parameters remain frozen. This significantly reduces the number of optimized parameters and enables efficient domain adaptation without sacrificing the general linguistic capabilities of the pretrained model. Specific LoRA configuration values are provided in Section IV.
@@ III-C Fine-Tuning on QA dataset
During this stage, all base model parameters remain frozen and only LoRA parameters are updated. Inputs are formatted explicitly as:
@@ IV-B1 Domain Corpus for DAPT
The heliophysics corpus is curated from peer-reviewed research articles published between 2010 and 2024, aligned with the operational era of the Solar Dynamics Observatory (SDO). Articles were selected from high-impact journals with an h5-index greater than 40 to ensure scientific rigor and relevance. These venues include The Astrophysical Journal (ApJ), The Astrophysical Journal Supplement Series (ApJS), Astronomy & Astrophysics (A&A), Space Weather, and Frontiers in Astronomy and Space Sciences, among others.
@@ IV-D Evaluation Method
We use win rate as the primary evaluation metric based on human pairwise comparison. For each test question, a human evaluator compares SolarGPT-QA’s response with a baseline model’s response and assigns one of three outcomes: Win, Tie, or Loss. The win rate is defined as the proportion of questions where SolarGPT-QA is judged to outperform the baseline. Results are reported by aggregating win, tie, and loss rates across all questions and separately across difficulty levels (Easy, Medium, Hard).
@@ IV-E Results
To further assess practical educational effectiveness, we conduct a small pilot student comprehension study involving 15 students. Each student was asked to read answers generated by each model and report whether they understood the response. Figure 6 shows the number of students who reported understanding the answers produced by each model.
ChatGPT (instruction-tuned, 10-shot) achieves the highest level of student comprehension, with all 15 students reporting understanding its responses. SolarGPT-QA follows closely, with 12 students indicating comprehension. ChatGPT (zero-shot) and Grok demonstrate moderate comprehension levels, while Claude, Gemini, and DeepSeek show lower student understanding overall. These results further suggests that instructional alignment and pedagogical optimization play a crucial role in educational question answering, beyond domain specialization alone.
@@ IV-F Ablation Study
We conduct an ablation study to analyze the impact of domain-adaptive pretraining (DAPT) and educational fine-tuning in SolarGPT-QA. We compare four model variants: the base LLaMA-3 model, LLaMA-3 with DAPT only, LLaMA-3 with QA fine-tuning only, and the full SolarGPT-QA model. DAPT significantly improves scientific grounding and correct use of heliophysics terminology but often produces explanations that are too technical for K–12 learners. QA fine-tuning alone improves readability and storytelling but lacks sufficient domain depth, leading to shallow or incomplete explanations. Table I shows SolarGPT-QA achieves the highest win rate (75%) and student comprehension (80%), suggesting that combining domain-adaptive pretraining with educational fine-tuning yields the strongest performance.
@@ V-A Discussion
The results presented in this work demonstrate that domain adaptation alone is insufficient for educational effectiveness, and that pedagogical alignment is a critical and often overlooked dimension when deploying large language models for science education. While general-purpose LLMs such as ChatGPT and Grok exhibit strong linguistic fluency and factual recall, our findings show that they frequently fail to deliver explanations that are accessible, coherent, and age-appropriate for K–12 learners in the context of space weather and heliophysics.
@@
These findings have broader implications beyond heliophysics. They suggest that many existing scientific LLMs—while impressive for expert-level analysis—are poorly suited for educational deployment without explicit adaptation. As AI systems increasingly enter classrooms, museums, and public outreach platforms, models that fail to account for learner expertise risk reinforcing misunderstanding or disengagement. SolarGPT-QA provides preliminary evidence that targeted domain adaptation combined with educational fine-tuning can effectively bridge this gap.
@@ V-B Future work
Second, we plan to introduce SolarGPT-PredEx, a multimodal prediction-and-explanation model that integrates time-series measurements and solar imagery with language modeling. Unlike the current SolarGPT-QA system, SolarGPT-PredEx will reason directly over observational data to produce both event predictions and data-grounded natural language explanations, enabling more transparent and interpretable space weather forecasting.
Third, future versions of SolarGPT-QA will incorporate retrieval-augmented generation (RAG), curriculum-aware prompting, and adaptive difficulty control to support classroom deployment and personalized learning. Larger and more diverse human evaluation studies—particularly involving teachers and students across multiple grade levels—will be conducted to rigorously assess educational effectiveness.
@@ V-C Conclusion
Human pairwise evaluation results show that while instruction-tuned ChatGPT (10-shot) achieves the strongest overall performance, SolarGPT-QA outperforms all zero-shot general-purpose baselines and suggests strong student comprehension and explanatory clarity. While these findings are preliminary and based on limited sample sizes, they provide initial evidence that integrating domain adaptation with educational fine-tuning can enhance the clarity and accessibility of scientific explanations. Ablation experiments further demonstrate that neither domain-adaptive pretraining nor educational fine-tuning alone is sufficient; rather, their combination is important for balancing scientific accuracy with pedagogical effectiveness.

### v2

@@ Abstract
Solar activity, including solar flares, coronal mass ejections (CMEs), and geomagnetic storms, can significantly impact satellites, aviation, power grids, data centers, and space missions. Extreme solar events can cause substantial economic damage with limited advance warning, underscoring the importance of early-warning systems, accurate forecasting, and effective education in space science. Although large language models (LLMs) perform well on general tasks, they often lack domain-specific knowledge and pedagogical capability to clearly explain complex space science concepts. We introduce SolarGPT-QA, a question answering system based on a domain-adapted large language model built on the LLaMA-3 base model. The model is trained using scientific literature and large-scale question-answer data generated with GPT-4 and refined using Grok-3 in a student-friendly storytelling style. Human pairwise evaluations show that SolarGPT-QA outperforms general-purpose models in zero-shot settings and achieves competitive performance compared to instruction-tuned models for educational explanations in space weather and heliophysics. A small pilot student comprehension study further suggests improved clarity and accessibility of the generated explanations. Ablation experiments indicate that combining domain-adaptive pretraining with pedagogical fine-tuning is important for balancing scientific accuracy and educational effectiveness. This work represents an initial step toward a broader SolarGPT framework for space science education and forecasting.
@@ I Introduction
Heliophysics provides a system-level understanding of the Sun and its influence throughout the heliosphere [37]. Within this broader heliophysical context, space weather science investigates the interconnected and multiscale interactions among the Sun, the solar wind, interplanetary space, and planetary environments, and examines how solar activity influences Earth’s space environment and technological systems [48]. Solar phenomena including solar flares (sudden bursts of electromagnetic radiation from the Sun) [48, 47], coronal mass ejections (CMEs) (large eruptions of magnetized plasma from the solar corona) [61], and solar energetic particles (SEPs) (high-energy charged particles accelerated by flares and CMEs) [45, 3] can trigger geomagnetic storms [21], that propagate through the near-Earth environment, disrupting satellite operations, communication networks, power grids, and data centers, while exposing astronauts and high-altitude aviation to elevated radiation hazards. For instance, the infamous ”Carrington Event” in 1859 [35] was the most intense geomagnetic storm on record, triggered by an extreme solar flare and CME, while the Quebec blackout in 1989 [16] vividly demonstrated the susceptibility of modern technological systems to space weather disturbances. Such disruptions have resulted in economic losses of billions of dollars, and an event of similar scale today could cause even greater societal and technological consequences in our technology-driven society [6, 3]. Additionally, as modern society becomes increasingly reliant on space-based technologies, including large satellite constellations such as Starlink [41], and as private space enterprises like SpaceX and Blue Origin pursue future space-based data centers and exploration missions [56], understanding space weather has emerged as a critical scientific, technological, and educational priority, with relevance extending as early as the high-school level, bringing an additional dimension to its importance.
@@
Large language models have become powerful tools for understanding and explaining complex information, with growing impact across many scientific fields [4]. They can answer questions, summarize research, and support learning. However, their use in space weather and heliophysics is still limited, largely because these areas rely on highly specialized literature and concepts that are rarely covered in general training data. One key challenge is that most space weather knowledge exclusively resides in technical papers and mission reports, which means general models often miss key terms, cause–effect links, or physical limits. Another issue is that teaching space science, particularly to K–12 students or early undergraduates, requires clear language, analogies, and a storytelling approach. General models are not built to balance accuracy that is tailored to the appropriate age.
@@
To fill this gap, we introduce SolarGPT-QA, a domain-adapted model designed for educational question answering in space weather and heliophysics for K–12 learners. The model is trained on a carefully curated corpus of high-impact, peer-reviewed research articles published since 2010, aligned with the Solar Dynamics Observatory (SDO) era of modern flare and CME research. These sources are further refined for relevance to solar eruptions and used to generate a large collection of synthetic question–answer pairs. Rather than targeting expert-level analysis, SolarGPT-QA focuses on helping students grasp core concepts through clear, accessible explanations to questions such as how solar flares form or why the Sun can disrupt technology on Earth.
@@ II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [20]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [30, 40, 22]. Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [60, 64], but their operational use is limited by computational cost and incomplete boundary conditions [63]. Although significant progress has been made in understanding solar magnetic field dynamics, a quantitative theory that rigorously connects the accumulation and restructuring of magnetic flux to the onset of extreme solar events remains lacking, leaving the key mechanisms governing flare and coronal mass ejection initiation still not fully understood [1]. This gap highlights the importance of data-driven approaches and machine learning methods for improving the prediction and characterization of extreme solar events.
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [23] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters. Progress in this area has been supported by the development of benchmark datasets. The Space Weather Analytics for Solar Flares (SWAN-SF) multivariate time series dataset integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [2]. Machine learning-based models in this domain include classical approaches such as logistic regression [50], support vector machines [7], and decision trees [33], as well as deep learning architectures for temporal sequence modeling, including LSTMs [36], contrastive learning approaches [57], graph-based methods [59], and transformer frameworks [58]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [18, 24].
Many efforts have applied deep learning on solar imagery, including magnetograms and extreme ultraviolet observations, to capture spatial precursors to flares and CMEs, while multimodal fusion of time series and images further enhances forecasting accuracy [39, 65, 25]. Despite these advances, most operational systems predict flares, CMEs, and SEPs independently and provide limited explanatory context, functioning largely as black boxes [15, 42, 43]. These limitations motivate unified, multimodal, and explainable forecasting frameworks that build on prior advances in time series modeling, dataset development, and interpretability.
@@ II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve state-of-the-art performance across wide range of the tasks and model scales [8, 14, 54, 55, 11, 10, 9]. There has been subsequent work and exploration for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [44, 67, 53, 46, 66]. DAPT and fine-tuning are common techniques to improve the LLM for specific datasets or domains. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [28, 26, 19]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [62, 38, 51, 13].
@@ III SolarGPT-QA
SolarGPT-QA is a domain-adapted large language model designed specifically for educational question answering in space weather and heliophysics, with an emphasis on K–12–friendly explanations. The methodology consists of four main stages: (1) domain corpus construction and preprocessing, (2) domain-adaptive pretraining (DAPT) using LoRA, (3) educational question–answer fine-tuning, and (4) inference-time constraints to promote clarity and reduce hallucinations. An overview of the system architecture is shown in Fig. 2. Let $\mathcal{D}_{H}$ denote a corpus of heliophysics literature, and $\mathcal{D}_{QA}$ denote a supervised question–answer dataset designed for K–12 learners. Our objective is to learn a conditional language model
@@
where $x$ is a heliophysics question, and $y$ is an age-appropriate explanatory answer that preserves scientific correctness while prioritizing conceptual clarity.
@@ III-A Domain Corpus Construction and Normalization
We curate a heliophysics corpus $\mathcal{D}_{H}=\{d_{i}\}_{i=1}^{N}$ from peer-reviewed literature published between 2010 and 2024, during the SDO operational period. Each document $d_{i}$ is extracted from PDF format and transformed using a deterministic preprocessing function:
@@ III-B Domain-Adaptive Pretraining with Low-Rank Adaptation
The overall training workflow is summarized in Algorithm 1. Only the low-rank matrices $A$ and $B$ are trainable, while the base model parameters remain frozen. This significantly reduces the number of optimized parameters, and enables efficient domain adaptation without sacrificing the general linguistic capabilities of the pretrained model. Specific LoRA configuration values are provided in Section IV.
@@ III-C Fine-Tuning on QA dataset
During this stage, all base model parameters remain frozen, and only LoRA parameters are updated. Inputs are formatted explicitly as:
@@ IV-B1 Domain Corpus for DAPT
The heliophysics corpus is curated from peer-reviewed research articles published between 2010 and 2024, corresponding to the operational era of the SDO. Articles were selected from high-impact journals with an h5-index greater than 40 to ensure scientific rigor and relevance. These venues include The Astrophysical Journal (ApJ), The Astrophysical Journal Supplement Series (ApJS), Astronomy & Astrophysics (A&A), Space Weather, and Frontiers in Astronomy and Space Sciences, among others.
@@ IV-D Evaluation Method
We use win rate as the primary evaluation metric based on human pairwise comparison. For each test question, a human evaluator compares SolarGPT-QA’s response with a baseline model’s response and assigns one of three outcomes: Win, Tie, or Loss. The win rate is defined as the proportion of questions where SolarGPT-QA is judged to outperform the baseline. Results are reported by aggregating win, tie, and loss rates across all questions, and separately across difficulty levels (i.e., Easy, Medium, Hard).
@@ IV-E Results
To further assess practical educational effectiveness, we conduct a small pilot student comprehension study involving 15 students. Each student was asked to read the answers generated by each model and report whether they understood the response. Figure 6 shows the number of students who reported understanding the answers produced by each model.
ChatGPT (instruction-tuned, 10-shot) achieves the highest level of student comprehension, with all 15 students reporting understanding its responses. SolarGPT-QA follows closely, with 12 students indicating comprehension. ChatGPT (zero-shot) and Grok demonstrate moderate comprehension levels, while Claude, Gemini, and DeepSeek show lower student understanding overall. These results further suggest that instructional alignment and pedagogical optimization play a crucial role in educational question answering, beyond domain specialization alone.
@@ IV-F Ablation Study
We conduct an ablation study to analyze the impact of domain-adaptive pretraining (DAPT) and educational fine-tuning in SolarGPT-QA. We compare four model variants: the base LLaMA-3 model, LLaMA-3 with DAPT only, LLaMA-3 with QA fine-tuning only, and the full SolarGPT-QA model. DAPT significantly improves scientific grounding and correct use of heliophysics terminology, but often produces explanations that are too technical for K–12 learners. QA fine-tuning alone improves readability and storytelling but lacks sufficient domain depth, leading to shallow or incomplete explanations. Table I shows SolarGPT-QA achieves the highest win rate (75%) and student comprehension (80%), suggesting that combining domain-adaptive pretraining with educational fine-tuning yields the strongest performance.
@@ V-A Discussion
The results presented in this work demonstrate that domain adaptation alone is insufficient for educational effectiveness and that pedagogical alignment is a critical and often overlooked dimension when deploying large language models for science education. While general-purpose LLMs such as ChatGPT and Grok exhibit strong linguistic fluency and factual recall, our findings show that they frequently fail to deliver explanations that are accessible, coherent, and age-appropriate for K–12 learners in the context of space weather and heliophysics.
@@
These findings have broader implications beyond heliophysics. They suggest that many existing scientific LLMs, while impressive for expert-level analysis, are poorly suited for educational deployment without explicit adaptation. As AI systems increasingly enter classrooms, museums, and public outreach platforms, models that fail to account for learner expertise risk reinforcing misunderstanding or disengagement. SolarGPT-QA provides preliminary evidence that targeted domain adaptation combined with educational fine-tuning can effectively bridge this gap.
@@ V-B Future work
Second, we plan to introduce SolarGPT-PredEx, a multimodal prediction-and-explanation model that integrates time-series measurements and solar imagery with language modeling. Unlike the current SolarGPT-QA system, SolarGPT-PredEx will directly reason over observational data to produce both event predictions and data-grounded natural language explanations, enabling more transparent and interpretable space weather forecasting.
Third, future versions of SolarGPT-QA will incorporate retrieval-augmented generation (RAG), curriculum-aware prompting, and adaptive difficulty control to support classroom deployment and personalized learning. Larger and more diverse human evaluation studies, particularly involving teachers and students across multiple grade levels, will be conducted to rigorously assess educational effectiveness.
@@ V-C Conclusion
Human pairwise evaluation results show that while instruction-tuned ChatGPT (10-shot) achieves the strongest overall performance, SolarGPT-QA outperforms all zero-shot general-purpose baselines, and suggests strong student comprehension and explanatory clarity. While these findings are preliminary and based on limited sample sizes, they provide initial evidence that integrating domain adaptation with educational fine-tuning can enhance the clarity and accessibility of scientific explanations. Ablation experiments further demonstrate that neither domain-adaptive pretraining nor educational fine-tuning alone is sufficient; rather, their combination is important for balancing scientific accuracy with pedagogical effectiveness.